OpenStudy (anonymous):

The value of x in the proportion 1/2:2/3 = 3/4:x is 1 4/9 1 7/9 1/4

OpenStudy (anonymous):

2/3 / 1/2 = 4/3 3/4 * 4/3 = 1 1